$SPAD/src/input fr1.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool fr1.output
)set message test on
)set message auto off
)clear all
--S 1 of 38
g := factor(4312)
--R
--R
--R         3 2
--R   (1)  2 7 11
--R                                                      Type: Factored(Integer)
--E 1
--S 2 of 38
unit(g)
--R
--R
--R   (2)  1
--R                                                        Type: PositiveInteger
--E 2
--S 3 of 38
numberOfFactors(g)
--R
--R
--R   (3)  3
--R                                                        Type: PositiveInteger
--E 4
--S 4 of 38
[nthFactor(g,i) for i in 1..numberOfFactors(g)]
--R
--R
--R   (4)  [2,7,11]
--R                                                          Type: List(Integer)
--E 4
--S 5 of 38

[nthExponent(g,i) for i in 1..numberOfFactors(g)]
--R
--R
--R   (5)  [3,2,1]
--R                                                          Type: List(Integer)
--E 5
--S 6 of 38
[nthFlag(g,i) for i in 1..numberOfFactors(g)]
--R
--R
--R   (6)  ["prime","prime","prime"]
--R                              Type: List(Union("nil","sqfr","irred","prime"))
--E 6
--S 7 of 38
factorList(g)
--R
--R
--R   (7)
--R   [[flg= "prime",fctr= 2,xpnt= 3], [flg= "prime",fctr= 7,xpnt= 2],
--R    [flg= "prime",fctr= 11,xpnt= 1]]
--IType: List(Record(flg: Union("nil","sqfr","irred","prime"),...
--E 7
--S 8 of 38
factors(g)
--R
--R
--R   (8)
--R   [[factor= 2,exponent= 3],[factor= 7,exponent= 2],[factor= 11,exponent= 1]]
--R                        Type: List(Record(factor: Integer,exponent: Integer))
--E 8
--S 9 of 38
first(%).factor
--R
--R
--R   (9)  2
--R                                                        Type: PositiveInteger
--E 9
)clear all
--S 10 of 38
g := factor(4312)
--R
--R
--R         3 2
--R   (1)  2 7 11

--R                                                      Type: Factored(Integer)
--E 10
--S 11 of 38
expand(g)
--R
--R
--R   (2)  4312
--R                                                        Type: PositiveInteger
--E 11
--S 12 of 38
reduce(*,[t.factor for t in factors(g)])
--R
--R
--R   (3)  154
--R                                                        Type: PositiveInteger
--E 12
)clear all
--S 13 of 38
g := factor(4312)
--R
--R
--R         3 2
--R   (1)  2 7 11
--R                                                      Type: Factored(Integer)
--E 13
--S 14 of 38
f := factor(246960)
--R
--R
--R         4 2   3
--R   (2)  2 3 5 7
--R                                                      Type: Factored(Integer)
--E 14
--S 15 of 38
f * g
--R
--R
--R         7 2   5
--R   (3)  2 3 5 7 11
--R                                                      Type: Factored(Integer)
--E 15
--S 16 of 38
f**500

--R
--R
--R         2000 1000 500 1500
--R   (4)  2    3    5   7
--R                                                      Type: Factored(Integer)
--E 16
--S 17 of 38
gcd(f,g)
--R
--R
--R         3 2
--R   (5)  2 7
--R                                                      Type: Factored(Integer)
--E 17
--S 18 of 38
lcm(f,g)
--R
--R
--R         4 2   3
--R   (6)  2 3 5 7 11
--R                                                      Type: Factored(Integer)
--E 18
--S 19 of 38
f + g
--R
--R
--R         3 2
--R   (7)  2 7 641
--R                                                      Type: Factored(Integer)
--E 19
--S 20 of 38
f - g
--R
--R
--R         3 2
--R   (8)  2 7 619
--R                                                      Type: Factored(Integer)
--E 20
--S 21 of 38
zero?(factor(0))
--R
--R
--R   (9)  true
--R                                                                Type: Boolean
--E 21

--S 22 of 38
zero?(g)
--R
--R
--R   (10)  false
--R                                                                Type: Boolean
--E 22
--S 23 of 38
one?(factor(1))
--R
--R
--R   (11)  true
--R                                                                Type: Boolean
--E 23
--S 24 of 38
one?(f)
--R
--R
--R   (12)  false
--R                                                                Type: Boolean
--E 24
--S 25 of 38
0$Factored(Integer)
--R
--R
--R   (13)  0
--R                                                      Type: Factored(Integer)
--E 25
--S 26 of 38
1$Factored(Integer)
--R
--R
--R   (14)  1
--R                                                      Type: Factored(Integer)
--E 26
)clear all
--S 27 of 38
nilFactor(24,2)
--R
--R
--R          2
--R   (1)  24
--R                                                      Type: Factored(Integer)

--E 27
--S 28 of 38
nthFlag(%,1)
--R
--R
--R   (2)  "nil"
--R                                                       Type: Union("nil",...)
--E 28
--S 29 of 38
sqfrFactor(30,2)
--R
--R
--R          2
--R   (3)  30
--R                                                      Type: Factored(Integer)
--E 29
--S 30 of 38
irreducibleFactor(13,10)
--R
--R
--R          10
--R   (4)  13
--R                                                      Type: Factored(Integer)
--E 30
--S 31 of 38
primeFactor(11,5)
--R
--R
--R          5
--R   (5)  11
--R                                                      Type: Factored(Integer)
--E 31
--S 32 of 38
h := factor(-720)
--R
--R
--R           4 2
--R   (6)  - 2 3 5
--R                                                      Type: Factored(Integer)
--E 32
--S 33 of 38
h - makeFR(unit(h),factorList(h))
--R
--R

--R   (7)  0
--R                                                      Type: Factored(Integer)
--E 33
)clear all
--S 34 of 38
p := (4*x*x-12*x+9)*y*y + (4*x*x-12*x+9)*y + 28*x*x - 84*x + 63
--R
--R
--R           2            2      2                  2
--R   (1)  (4x  - 12x + 9)y  + (4x  - 12x + 9)y + 28x  - 84x + 63
--R                                                    Type: Polynomial(Integer)
--E 34
--S 35 of 38
fp := factor(p)
--R
--R
--R                2  2
--R   (2)  (2x - 3) (y  + y + 7)
--R                                          Type: Factored(Polynomial(Integer))
--E 35
--S 36 of 38
D(p,x)
--R
--R
--R                  2
--R   (3)  (8x - 12)y  + (8x - 12)y + 56x - 84
--R                                                    Type: Polynomial(Integer)
--E 36
--S 37 of 38
D(fp,x)
--R
--R
--R                   2
--R   (4)  4(2x - 3)(y  + y + 7)
--R                                          Type: Factored(Polynomial(Integer))
--E 37
--S 38 of 38
numberOfFactors(%)
--R
--R
--R   (5)  3
--R                                                        Type: PositiveInteger
--E 38
)spool

)lisp (bye) ———-

10

References^{[1] nothing}

11